OpenStudy (anonymous):

The hypotenuse of a right triangle is 6 inches longer than the shortest side and 3 inches longer than the remaining side. Find the dimensions of the triangle

OpenStudy (anonymous):

Let c be the hypotenuse (c-6)^2+(c-3)^2=c^2 c=3 or c = 15 c=3 doesn't make sense in this situation so the hypotenuse = 15, shortest side=9, the remaining side=12

OpenStudy (anonymous):

TY